# Journal Title

### v1

@@ 4 Perspectives
Lastly we discuss the scalability of the described numerical methods from simulating a reduced water experiment to the industrial steel-slag-argon system. From a modelling perspective, the principal difference between the used mathematical model and the physical system is that it is known that the compressibility of air can no longer be ignored [15, 28]. This could affect the dynamics in the bubble plume, though its effect on macroscopic quantities such as the open-eye size and species mass flux is uncertain. More importantly, the Reynolds number of the industrial system increases from $10^{4}$ to $10^{6}$ when compared to the reduced system. Using the scaling argument of Pope [40], which says that the computational cost of a turbulent system scales with $Re^{11/4}$, estimates that simulating the industrial system would cost roughly 300 thousand times more than the reduced system. Using AMR generally adjusts the coefficient rather than exponent and it can be assumed the AMR cell reduction in the model system would be of the order of that seen in the industrial system. As the reduced scale simulations are already expensive, ($\sim 10^{5}\mathrm{\;to}\;10^{6}$ CPU-hours), simulating the industrial system is beyond the computational feasibility required for the current approach to be integrated into the engineering design cycle. To combat this, one could introduce LES-based turbulence modelling, which has been successfully applied to simulate similar Reynolds number flows at moderate costs [41]. Furthermore the morphological complexity of the system could be simplified by using morphological-multiscale approaches. The simplest example could be to replace the interface-resolved bubble plume with an Euler-Euler model, using a tracer to represent the local buoyancy and a bubble turbulent Schmidt number to simulate the dispersion of bubbles. A more challenging improvement would be to use a combination of interface-resolved methods and Euler-Lagrange approaches to resolve the large interfaces and use Lagrangian particles to model the small detached droplets of the slag layer [42]. The mass flux from the Lagrangian particles would be computed using empirical correlations. Lastly switching to more advanced high performance computing technologies (e.g. GPU-accelerated computing) could save roughly one order of magnitude on energy consumption [43]. Importantly, all of the described improvements are existing technologies which are compatible with the methods used in this paper.

### v2

@@ 4.2 Scalability of the model
As a last remark, we discuss the scalability of the described numerical methods from simulating a reduced water experiment to the industrial steel-slag-argon system. From a modeling perspective, the principal difference between the used mathematical model and the physical system is that it is known that the compressibility of air can no longer be ignored [17, 34]. This could affect the dynamics in the bubble plume, though its effect on macroscopic quantities such as the open-eye size and species mass flux is uncertain. More importantly, the Reynolds number of the industrial system increases from $10^{4}$ to $10^{6}$ when compared to the reduced system. Using the scaling argument of Pope [48], which says that the computational cost of a turbulent system scales with $Re^{11/4}$, estimates that simulating the industrial system would cost roughly 300 thousand times more than the reduced system. Using adaptive mesh resolution (AMR) generally adjusts the coefficient rather than the exponent and it can be assumed the AMR cell reduction in the model system would be of the order of that seen in the industrial system. As the reduced scale simulations are already expensive, ($\sim 10^{5}\mathrm{\;to}\;10^{6}$ CPU-hours), simulating the industrial system is beyond the computational feasibility required for the current approach to be integrated into the engineering design cycle. To combat this, one could introduce LES-based turbulence modelling, which has been successfully applied to simulate similar Reynolds number flows at moderate costs [49]. Furthermore the morphological complexity of the system could be simplified by using morphological-multiscale approaches. The simplest example could be to replace the interface-resolved bubble plume with an Euler-Euler model, using a tracer to represent the local buoyancy and a bubble turbulent Schmidt number to simulate the dispersion of bubbles. A more challenging improvement would be to use a combination of interface-resolved methods and Euler-Lagrange approaches to resolve the large interfaces and use Lagrangian particles to model the small detached droplets of the slag layer [50]. The mass flux from the Lagrangian particles would be computed using empirical correlations. Lastly switching to more advanced high performance computing technologies (e.g. GPU-accelerated computing) could save roughly one order of magnitude on energy consumption [51]. Importantly, all of the described improvements are existing technologies which are compatible with the methods used in this paper.
5 Conclusions
In this paper we have investigated hydrodynamics and mass transfer in two reduced-scale model metallurgical ladles by means of direct numerical simulation (DNS), comparing our results with experimental data from the literature. The numerical method presented shows a significant advancement on the state of the art for computational mass transfer analysis in reduced-scale metallurgical ladles. The results shown demonstrate that combining multiscale approaches with adaptive mesh resolution to treat ranges of length scales in the hydrodynamic system together with a subgrid-scale (SGS) model to treat the thin concentration boundary layers, yields relative agreement between experiments and simulations. To the best of the authors knowledge, this is achieved for the first time without extrapolating results simulated with artificially high diffusivities, as the SGS model allowed us to simulate very large $Sc$ numbers with computational costs comparable to simulating just hydrodynamics without mass transfer.
Having good agreement between simulations and experiments is crucial, as the former can give much richer data than what is measurable in experiments. For example, we can directly measure the time-averaged velocity field to understand physical phenomena present in the system rather than rely on macroscopic, optically-measurable quantities such as the open-eye size. We can also directly measure the local distribution mass flux at the oil-water interface, which gives a better understanding of the relationship between the input flow rate and the output flux. We can use this rich data to inform a phenomenological model that predicts the measured quantities as a function of the flow rate. A mechanistic model that explains the measured quantities and attempts to unify all hydrodynamic and mass transfer experiments is under development and will be the subject of a future paper.
@@ 5 Conclusions
Appendix A Structure of the bubble plume considering bouyant and momentum driven effects
We consider the relative effects of bouyancy and momentum-driven effects on the structure of the bubble plume. We assume the classical scaling for the Boussinesq buoyant plume with the rise velocity given by [52, 53]
${\overline{u}_{z}}(z)\sim\left(\frac{gQ}{z}\right)^{1/3}.$ (30)
For a self-similar free jet, the velocity follows [52, 53]
${\overline{u}_{z}}(z)\sim\frac{(M/\rho_{w})^{1/2}}{z}$ (31)
where the momentum flux $M$ is given by
$M=\int_{S_{z}}\langle\rho u^{2}_{z}\rangle{\mathrm{d}}S$ (32)
Assuming a plug flow in the inlet, the momentum flux may be estimated at the inlet by $M=\rho_{a}\langle V_{I}^{2}\rangle\pi d_{I}^{2}/4$ where $V_{I}$ is the inlet velocity, related to the flow rate by $Q/(4\,d_{i}^{2})$, leading to relation between momentum flux and volume flux
$M=\frac{4\rho_{a}}{\pi d_{I}^{2}}Q^{2}$ (33)
and the velocity scaling
${\overline{u}_{z}}(z)\sim\left(\frac{4\rho_{a}}{\pi\rho_{w}}\right)^{1/2}\frac{Q}{d_{I}z}$ (34)
A rough approximation may obtained by combining the two scalings
${\overline{u}_{z}}(z)\simeq c_{u,z}\left(\frac{gQ}{z}\right)^{1/3}+c_{u,z,m}\left(\frac{4\rho_{a}}{\pi\rho_{w}}\right)^{1/2}\frac{Q}{d_{I}z},$ (35)
where we have introduced constants $c_{u,z}$ and $c_{u,z,m}$. The ratio of the plume (buoyancy-driven) velocity and of the free-jet (momentum-driven) velocity, that is the ratio of the two terms of Eq. (35) at height of the top of the water layer $h_{w}$ is
$R=\frac{h_{w}}{d_{I}}\left(\frac{Q^{2}}{gh_{w}^{5}}\right)^{1/3}\left(\frac{4\rho_{a}}{\pi\rho_{w}}\right)^{1/2}$ (36)
For all flow rates chosen in the two test-cases, the value computed by Eq. 36 is small, showing momentum-driven effects are negiligble on the development of the structure of the bubble plume.
Appendix B Relationship between the oil-water interface area and mass transfer
We consider the relationship between the oil-water exchange area and mass transfer. This follows the principle of the interfacial area concentration, a geometric quantity commonly used in mixture models of bubbly flows to characterize mass, momentum and energy transfer, usually characterized as the interfacial area per unit volume [54]. We assume that the height of the water layer has little effect on the total oil-water interfacial area, so that instead we define an order one non-dimensional quantity $\mathrm{A_{int}^{*}}$, given by the total oil-water interfacial area divided by the cross sectional area of the tank evaluated at the height of the oil layer. We plot both instantaneous and time-averaged values of the non-dimensional interfacial area in Fig. 12 and 13. We observe that in the Kim & Fruehan case, the interfacial area only increases slightly at the lowest flow rates and then increases dramatically for the highest flow rate, corresponding to the atomising regime. More strikingly in the Joubert case, the open-area decreases with increasing flow rates in the non-atomising regime before increasing when the oil layer atomises. To explain this phenomenon, we consider a simplified approximation of the oil-water surface area without considering fragmentation
$\mathrm{A_{int}}=\mathrm{A_{cr}}-\pi r_{\mathrm{oe}}^{2}+2\pi r_{\mathrm{oe}}h,$ (37)
where $\mathrm{A_{cr}}$ is the cross sectional area, $r_{\mathrm{oe}}$ is the open-eye radius and $h$ is the height of the oil layer given by
$h=\frac{V_{o}}{\mathrm{A_{cr}}-\pi r_{\mathrm{oe}}^{2}},$ (38)
where $V_{o}$ is the oil volume. As the open-eye radius increases, the total exchange area would thus decrease. This phenomenon is not quite observed in the case of the Kim & Fruehan case, which is caused both by the open-eye area being larger relative to the tank cross-sectional area, consequently causing a less significant effect, and by the larger interface deformation and smaller amounts of oil fragmentation observed at intermediate flow rates for these simulations.
Since the evolution of the interfacial area in the non-atomising regime does not correspond to the observed increase in mass transfer rate, it is demonstrated that the interfacial area does not correlate very well with the mass transfer rate in the ladle experiments, at least for low to intermediate flow rates. This is different to what is commonly observed for mass transfer in turbulent bubbly flows, where it can often be assumed that the Sherwood number of each bubble is either statistically uniform or given by a known distribution. Consequently, the global Sherwood number would be the product of the mean bubble Sherwood number and the interfacial area.
In the ladle case, we can consider three topologically distinct regions: the annulus around the open-eye, the relatively flat region between the oil and water bulks outside the annulus, and all the detached droplets. Each region is governed by different flow and mass transfer phenomena, namely forced convection in a shear layer for the annulus, free convection for the outer region, and droplet mediated mass transfer for the droplets. It is observed that the Sherwood number in the annulus region has a dominant role in the overall mass transfer rate, at least for the flow rates considered. It is expected that at higher flow rates than the ones simulated, mass transfer by fragmented droplets would play a more significant role, thus corresponding to the shift in mass transfer regimes observed at high flow rates in experiments.
Figure 12: Instantaneous nondimensional oil-water interfacial area computed in the quasi-steady regime for (a) the Kim & Fruehan simulations and (b) the Joubert simulations
Figure 13: Time-averaged nondimensional oil-water interfacial area for both sets of simulations. The standard error is used as error bars.